OpenStudy (anonymous):

64x12−−−−−√4y14 simplify by factoring assume that all expressions under the redical represent nonegative numbers needs an exact anser use radicals if needed

OpenStudy (anonymous):

I don't fully understand the equation given, is it: \[64 * 12 = \sqrt{4y} * 14\]

OpenStudy (anonymous):

it got messed up when i tried to copy it

OpenStudy (anonymous):

\[\sqrt[4]{64x ^{12}}y ^{14}\]

OpenStudy (anonymous):

Now that makes sense. You can start by converting the root to a power (a root is just a fractional power): \[(64x ^{12})^{1/4}*y^{14}\] Then distribute the power: \[\sqrt[4]{64}*x^3*y^{14}\] And simplify: \[2\sqrt{2}x^3y^{14}\]

OpenStudy (anonymous):

says it is wrong according to the study guide

OpenStudy (anonymous):

says needs an exact answer and to use radicals if necessary

OpenStudy (anonymous):

Is this online? Do you have the answer (if so, give it and I'll find my error faster)? Make sure you are entering it correctly (not including x in the radical, try using * to multiply the variables, etc). Online homework can be rather picky sometimes. I'll re-evaluate a while.

OpenStudy (anonymous):

Try using (x^12)^(1/4) instead of x^3. Using x^3 assumes x is positive.

OpenStudy (anonymous):

\[2x ^{3}y ^{3}\sqrt[4]{4y ^{2}}\]

OpenStudy (anonymous):

Ah, the y^{14} was under the radical as well (not shown in your response)...\[\sqrt[4]{64x^{12}y^{14}}=\sqrt[4]{4*16*x^{12}*y^2*y^{12}}\] You can take out the 16, x^12, and y^12 of the radical, leaving 4y^2 in the radical.

OpenStudy (anonymous):

the answer is \[4x^{3}y^{3}\sqrt[4]{y^{2}}\]